OpenStudy (monkeyagf):

Which of the following is the solution to the equation 3 over 4y = 24? A. y = 18 B. y = 32 C. y = 64 D. y = 72

OpenStudy (monkeyagf):

@AloneS

alones (alones):

3/4y =24 It -0should look like this \(\frac{ 3 }{ 4y } \) =24 multiply both sides by \(\frac{ 4 }{ 3 }\) \(\frac{ 4 }{ 3 }\) * \( \frac{3 }{ 4y }\) = \( \frac{ 4 }{ 3 }\) *24 y=32

OpenStudy (benthicflare):

AloneS, I think you are saying \[\frac{ 3 }{ 4 } y=24\] Multiplying \[\frac{ 4 }{ 3 }\] on both sides solves the problem

OpenStudy (monkeyagf):

ok, thx